Suppose 𝑓 and 𝑔 are continuous functions on [π‘Ž, 𝑏] and that

𝑔(π‘₯) β‰₯ 0 for all π‘₯ ∈ [π‘Ž, 𝑏]. Prove that there exists π‘₯ in [π‘Ž, 𝑏] such that

𝑏 𝑏

∫ 𝑓(𝑑)𝑔(𝑑)𝑑𝑑 = 𝑓(π‘₯) ∫ 𝑔(𝑑)𝑑𝑑.

π‘Ž π‘Ž



1
Expert's answer
2021-06-16T10:26:44-0400

Proof.

Since "f" is continuous on the segment "[a,b]" then "f" is bounded. We denote "M=\max _{ [a,b] }{ f(x)\ ,\ m=\min _{ [a,b] }{ f(x) } }" .Thus "m\le f(x)\le M" . Multiplying by "g(x)" , we obtain the inequality "mg(x)\le f(x)g(x)\le Mg(x)" . All function in the inequality are continuous , hence integrable. We integrate the inequality:

"m\int _{ a }^{ b }{ g(x)dx\le \int _{ a }^{ b }{ f(x)g(x)dx\le M\int _{ a }^{ b }{ g(x)dx } } }" (1)

Since, "g(x)\ge 0" , then "\int _{ a }^{ b }{ g(x)dx } \ge 0" . If "\int _{ a }^{ b }{ g(x)dx } =0" , then from (1) it follows "\int _{ a }^{ b }{ f(x)g(x)dx } =0" . In this case the equality

"\int _{ a }^{ b }{ f(x)g(x)dx } =f(x)\int _{ a }^{ b }{ \ g(x)dx }" (2)

has the form "0=f(x)\cdot 0" and is satisfied for any "x\in \left[ a,b \right]" .

If "\int _{ a }^{ b }{ \ g(x)dx } \neq 0\quad", then "\int _{ a }^{ b }{ \ g(x)dx } >0\quad". Denote "c=\frac { \int _{ a }^{ b }{ f(x)g(x)dx } }{ \int _{ a }^{ b }{ \ g(x)dx } }" .

Dividing all parts of inequality (1) by "\int _{ a }^{ b }{ \ g(x)dx }" we have "m\le c\le M" . By the Theorem on the intermediate value of a continuous function there is a point "x\in \left[ a,b \right]" such that "f(x)=c." So there exists point "x\in \left[ a,b \right]" such that equality (2) is satisfied.